# references.bib

### v1

@@ 4.1 Wasserstein gradient flows in the case $\lambda>0$
Let the assumptions of Lemma 3.1 hold and let $d\geq 3$. Consider some initialization $\mu_{0}\in\mathcal{P}(\Omega)$ and, for some $N\geq 0$, note $\hat{\mu}_{0}\coloneqq N^{-1}\sum_{i=1}^{N}\delta_{\omega_{i}}$ the empirical measure where $\{\omega_{i}\}_{1\leq i\leq N}$ are i.i.d. samples of $\mu_{0}$. For $\lambda>0$, let $(\mu^{\lambda}_{t})_{t\geq 0}$ and $(\hat{\mu}_{t}^{\lambda})_{t\geq 0}$ be the gradient flow of $\mathcal{L}^{\lambda}_{f}$ starting from $\mu_{0}$ and $\hat{\mu}_{0}$ respectively. Then there exists a constant $A=A(d,\Omega)$ s.t. for every $t\geq 0$ and every $\varepsilon>0$:
$\mathbb{P}\left(\mathcal{W}_{1}(\hat{\mu}^{\lambda}_{t},\mu^{\lambda}_{t})\geq\varepsilon)\right)\leq\frac{A}{\varepsilon}N^{-1/d}e^{Ct}\,,$
where $C=C(f^{*},\lambda)$ is the constant in Theorem 4.1.
@@
for some constant $A=A(d,\Omega)>0$ depending on the dimension and on the domain $\Omega$. Then using the contraction rate in Theorem 4.1 we have a constant $C=C(f^{*},\lambda)>0$ such that for every $t\geq 0$:
@@
The result the follows by applying Markov’s inequality. ∎
@@ 4.2 Wasserstein gradient flows in the case $\lambda=0$ and ultra-fast diffusions
$\partial_{t}\mu_{t}=-\|\bar{\nu}\|^{r}\mathrm{div}\left(\mu_{t}\nabla\left(\frac{\bar{\mu}}{\mu_{t}}\right)^{r}\right).$ (37)
@@
$\partial_{t}\mu_{t}=-r\|\bar{\nu}\|^{r}_{\mathrm{TV}}\mathrm{div}\left(\bar{\mu}\nabla\left(\frac{\bar{\mu}}{\mu_{t}}\right)^{r-1}\right)$
@@
Definition 4.2 (Gradient flow of $\mathcal{L}^{0}_{r}$ (Def. 1.1 in [iacobelli2019weighted])).
Let $\mu_{0}\in\mathcal{P}(\Omega)$ admit a density $\mu_{0}\in L^{r+2}(\Omega)$. We say $(\mu_{t})_{t\geq 0}$ is a weak solution of eq. 37 or a gradient flow for $\mathcal{L}^{0}_{r}$ starting from $\mu_{0}$ if it a narrowly continuous curve in $\mathcal{P}(\Omega)$ with $\lim_{t\to 0^{+}}\mu_{t}=\mu_{0}$, s.t.
@@
$\left(\frac{\mu_{t}}{\bar{\mu}}\right)^{r-1}\in L^{2}_{loc}((0,\infty),H^{1}(\Omega)),\quad\frac{\bar{\mu}}{\mu_{t}}\in L^{2}_{loc}((0,\infty),H^{1}(\Omega))\,.$
@@
In [iacobelli2019weighted], the authors establish the existence of uniqueness of gradient flows for the functional $\mathcal{L}^{0}_{r}$. More precisely, they show that, under appropriate assumptions on the initialization $\mu_{0}$ and on the target $\bar{\mu}$, the iterates of the proximal scheme in eq. 30 converge towards a curve $(\mu_{t})_{t\geq 0}$ that is a gradient flow of the functional $\mathcal{L}^{0}_{r}$ in the sense of Definition 4.2.
Theorem 4.2 (Well-posedness of the gradient flow equation for $\lambda=0$ (Thm. 1.2 of [iacobelli2019weighted])).
Assume $\mu_{0}$ and $\bar{\mu}$ are absolutely continuous and have bounded log-densities. Then there exists a unique weak solution of eq. 37 starting from $\mu_{0}$ in the sense of Definition 4.2.
@@
Theorem 4.3 (Theorem 1.4 in [iacobelli2019weighted]).
Assume $\mu_{0}$ and $\bar{\mu}$ are absolutely continuous and have bounded log-densities. For $\mu_{0}\in\mathcal{P}(\Omega)$, let $(\mu_{t})_{t\geq 0}$ be a weak solution of eq. 37 starting from $\mu_{0}$ in the sense of Definition 4.2. Then in fact the log-density of $\mu_{t}$ is bounded uniformly over $t\geq 0$ and there exists a constant $C=C(\Omega,\bar{\mu},\mu_{0})>0$ s.t. for every $t\geq 0$:
@@
For completeness, we give here some of the key arguments of the proof of the above Theorem 4.3 in the case where $r=2$. In this case we have for every $\mu\in\mathcal{P}(\Omega)$:
@@
where $\chi^{2}$ is the chi-square divergence. The following Lemma 4.3 establishes the desired linear convergence rate for the proximal scheme defined in eq. 30 with the loss $\mathcal{L}^{0}_{2}$. The result in continuous time then follows then follows from the lower semi-continuity of the $\chi^{2}$-divergence as the curve $(\mu_{t})_{t\geq 0}$ is obtained by taking the limit of the discrete process $(\mu^{\tau}_{k})_{k\geq 0}$ when the discretization time $\tau$ tends to zero.
From a technical perspective, the proof of Lemma 4.3 relies on a Poincaré inequality satisfied by $\mu_{t}$. It is indeed well-known that such inequality controls the convergence rate of Fokker-Planck equations towards their stationary distribution distribution in $\chi^{2}$-distance [pavliotis2014stochastic, Thm. 4.4]. This can for example be used to prove the convergence of sampling algorithms such as Langevin Monte Carlo [chewi2024analysis, chewi2020svgd]. In our case, the ultra-fast diffusion eq. 37 is to be interpreted as a Wasserstein gradient flow for $\mathcal{L}^{0}_{2}$, which, by the above eq. 39, is the reverse $\chi^{2}$-divergence between $\mu_{t}$ and $\bar{\mu}$ and the convergence rate is controlled by the Poincaré constant of $\mu_{t}$. This rate may thus a priori evolve and vanish during training but, crucially, [iacobelli2019weighted, Lem. 2.4] shows that it is here a property of solutions to the ultra-fast diffusion equation that the log-density ratio $\|\log\left(\frac{\bar{\mu}}{\mu_{t}}\right)\|_{\infty}$ decreases with time. As a consequence, it is sufficient to assume that the log-density is bounded at initialization to obtain a control over the Poincaré constant of $\mu_{t}$, for $t\geq 0$, by a classical perturbation argument [ane2000inegalites, Thm. 3.4.1].
@@
Remark 4.2 (Dependence of the convergence rate w.r.t. dimension).
It follows from the proof that the convergence rate $C$ in Lemma 4.3 scales linearly with $C_{P}(\Omega)^{-1}$ where $C_{P}(\Omega)$ is the Poincaré constant of the domain $\Omega$. For bounded, Lipschitz and convex domains of $\mathbb{R}^{d}$ or for the flat Torus $\mathbb{T}^{d}$, this constant is in particular independent of the dimension $d$ [payne1960optimal]. Therefore, the correspondence between the training of neural networks in the two-timescale regime and solutions to ultra-fast diffusions points towards the fact that gradient methods, with suitable hyperparameter scaling, are amenable to efficient feature learning in the training of neural networks, without suffering from the curse of dimensionality [donoho2000high]. Note however that the convergence rate $C$ in Lemma 4.3 is exponentially bad in the log-density ratio $\|\log(\bar{\mu}/\mu_{0})\|_{\infty}$. In particular the convergence rate does not hold in case the teacher feature distribution is supported on a finite number of atoms.
@@ 5 Convergence of gradient flow
The main purpose of this work is to study in what extend the gradient flow dynamics defined Definitions 4.1 and 4.2 allow to recover the teacher feature distribution $\bar{\mu}$ associated to the observed signal $Y$ in Assumption 1. Whereas Theorem 4.3 shows convergence of the gradient flow of $\mathcal{L}^{0}_{f}$, that is solutions to the ultra-fast diffusion eq. 37, such dynamics are hardly numerically tractable in practice due to the absence of the regularization parameter $\lambda$. For this reason we are interested here in the asymptotic behavior of the gradient flow of $\mathcal{L}^{\lambda}_{f}$ in the case where $\lambda>0$. A difficulty is that, in the case $\lambda=0$, the proof of Theorem 4.3 relies on the implicit behavior of eq. 37 which preserves the density of solutions. Such a behavior is a priori not expected to hold when $\lambda>0$. As a consequence, the result of this section holds under supplementary regularity assumptions on the solutions to eq. 33.
@@ 5.1 Algebraic convergence rate
At fixed $\lambda>0$, we are able to obtain convergence towards the minimizer $\bar{\mu}^{\lambda}_{f}$ of $\mathcal{L}^{\lambda}_{f}$ under mild regularity assumptions on solutions to the gradient flow eq. 33. Specifically, for a probability measure $\mu\in\mathcal{P}(\Omega)$ and a function $h\in\mathcal{C}^{1}$ we define the weighted Sobolev semi-norm of $h$ as:
@@
Then for a measure $\nu\in\mathcal{M}(\Omega)$ s.t. $\int_{\Omega}\mathrm{d}\nu=0$ the negative weighted Sobolev semi-norm $\|\nu\|_{\dot{H}^{-1}(\mu)}$ is defined by duality as:
@@
The following Theorem 5.1 states convergence of $\mathcal{L}^{\lambda}_{f}$ towards 00 with an algebraic convergence rate provided $\|\mu_{t}-\frac{\bar{\nu}}{\bar{m}}\|_{\dot{H}^{-1}(\mu_{t})}$ stays bounded along the gradient flow. As discussed below, since the domain $\Omega$ is compact, this assumption is satisfied for example when both distribution have bounded log-densities. The arguments are similar to the one presented in [glaser2021kale], where the authors consider an infimal convolution between a kernel discrepancy and the $\mathrm{KL}$-divergence. Importantly, the obtained convergence rate depends on the bound on $\|\mu_{t}-\frac{\bar{\nu}}{\bar{m}}\|_{\dot{H}^{-1}(\mu_{t})}$ but is independent of $\lambda>0$.
Theorem 5.1.
@@
Note that it follows from the assumptions on $f$ and $\bar{\nu}$ that $\inf\mathcal{L}^{\lambda}_{f}=0$ and that this infimum is attained only for $\mu=\bar{\nu}/\bar{m}$. Thus, the last statement on the convergence of $\mu_{t}$ follows from the convergence of $\mathcal{L}^{\lambda}_{f}(\mu_{t})$ to 00 and from the lower-semicontinuity of $\mathcal{L}^{\lambda}_{f}$ (see section 3).
To obtain the convergence rate, consider $\mu\in\mathcal{P}(\Omega)$ and note that by eq. 19 we have $\mathcal{L}^{\lambda}_{f}(\mu)=\max_{\alpha}\mathcal{K}(\alpha,\mu)$ where for every $\alpha\in L^{2}(\rho)$:
@@
where $f^{*}$ is the Legendre transform of $f$. Let us note $\alpha^{\lambda}=\alpha^{\lambda}_{f}[\mu]$ the maximizer of $\mathcal{K}(\alpha,\mu)$. Then, using the convexity of $f^{*}$ we have for every $\omega\in\Omega$:
@@
Also by assumptions $\partial f(\bar{m})=0$ and hence by properties of the Legendre transform $\partial f^{*}(0)=\bar{m}$. Also $f^{*}(0)=-f(\bar{m})=0$ and after integrating w.r.t. $\bar{\nu}$ :
@@
Let us comment on the assumptions of Theorem 5.1. The second assumption specifically, is automatically satisfied in case $\bar{\nu}$ has bounded density and $\mu_{t}$ has bounded log-density, uniformly over $t\geq 0$. Indeed, for $\mu\in\mathcal{P}(\Omega)$ having a lower-bounded log-density we have that the weighted Sobolev semi-norm .$\|.\|_{\dot{H}^{1}(\mu)}$ is lower-bounded by the classical Sobolev semi-norm .$\|.\|_{\dot{H}^{1}(\pi)}$ where we recall that $\pi$ is the (normalized) Lebesgue measure over $\Omega$. Precisely, if $\pi\ll\mu$ and $\mathrm{d}\pi/\mathrm{d}\mu\leq C_{1}$ then for every $f\in\mathcal{C}^{1}$:
@@
In this case, the weighted negative Sobolev semi-norm .$\|.\|_{\dot{H}^{-1}(\mu)}$ is upper-bounded by the semi-norm .$\|.\|_{\dot{H}^{-1}(\pi)}$ and for every $\nu\in\mathcal{M}(\Omega)$ with $\int_{\Omega}\mathrm{d}\nu=0$ we have:
@@
Moreover, this last quantity can be estimated by the Wasserstein distance. Indeed, for probability measures having bounded log-densities, the Wasserstein distance $\mathcal{W}_{2}$ is equivalent to the negative Sobolev semi-norm .$\|.\|_{\dot{H}^{-1}(\pi)}$. If $\mu,\nu\in\mathcal{P}(\Omega)$ are such that $\frac{\mathrm{d}\mu}{\mathrm{d}\pi},\frac{\mathrm{d}\nu}{\mathrm{d}\pi}\leq C_{2}$ for some constant $C_{2}>0$ we have [santambrogio2015optimal, Lem. 5.33 and Thm. 5.34]:
@@
Finally, the Wasserstein distance $\mathcal{W}_{2}(\mu,\nu)$ is always bounded by $\mathrm{diam}(\Omega)$ which is finite, hence ensuring the second assumption of Theorem 5.1 is satisfied.
@@
The algebraic convergence rate stated in the above Theorem 5.1 in the case $\lambda>0$ stands in contrast with the faster linear convergence stated in Theorem 4.3 in the case $\lambda=0$. For this reason, we are interested in comparing the gradient flow dynamics with and without regularization.
Below we assume $f(t)=|t|^{r}/(r-1)$ for some $r>1$ and Theorem 5.2 shows local uniform in time convergence of gradient flows of $\mathcal{L}^{\lambda}_{r}$ to gradient flows of $\mathcal{L}^{0}_{r}$, i.e. solutions to the ultra-fast diffusion equation eq. 36, when the regularization strength $\lambda$ vanishes. To obtain such a result we assume regularity on the density ratio $\frac{\mathrm{d}\bar{\nu}}{\mathrm{d}\mu^{\lambda}_{t}}$. Namely, we assume that the Legendre-conjugate $\partial f(\frac{\mathrm{d}\bar{\nu}}{\mathrm{d}\mu^{\lambda}_{t}})$ stays bounded in the RKHS $\mathcal{H}$, defined as the image of the convolution operator $\Phi^{\top}:L^{2}(\rho)\to\mathcal{C}^{0}(\Omega)$ (eq. 41). Using classical results from the theory of inverse problems, such a source condition ensures the dual variable $\alpha\in L^{2}(\rho)$ stays uniformly bounded for $\lambda>0$ (Lemma 5.1). Therefore, provided $\mathcal{H}$ is sufficiently regular, such a regularity assumption ensures compactness of the Wasserstein gradient $\nabla\mathcal{L}^{\lambda}_{r}[\mu_{t}]=\nabla f^{*}(\Phi^{\top}\alpha^{\lambda}_{t})$ in $\mathcal{C}^{1}$ and allows passing to the limit in eq. 34 to obtain eq. 38.
Theorem 5.2.
Assume Assumption 1 hold with $\bar{\nu}$ a positive measure with bounded log-density, $f(t)=|t|^{r}/(r-1)$ for some $r>1$ and the assumptions of Theorem 4.1 and Theorem 4.2 are satisfied. Consider some initialization $\mu_{0}\in\mathcal{P}(\Omega)$ s.t. $\mu_{0}$ has bounded log-density. For $\lambda\geq 0$ let $(\mu^{\lambda}_{t})_{t\geq 0}$ be the gradient flow of $\mathcal{L}^{\lambda}_{r}$, starting from $\mu_{0}$ in the sense of Definition 4.1 (when $\lambda>0$) and Definition 4.2 (when $\lambda=0$). Moreover, assume $\mathcal{H}$ is compactly embedded in $\mathcal{C}^{1}(\Omega)$ and $\partial f(\frac{\mathrm{d}\bar{\nu}}{\mathrm{d}\mu^{\lambda}_{t}})$ is bounded in $\mathcal{H}$, locally uniformly over $t\geq 0$ and uniformly over $\lambda>0$. Then for any $T\geq 0$:
@@
The curve $(\mu_{t})^{\lambda}$ for $\lambda>0$ are gradient flows for the functionals $\mathcal{L}^{\lambda}_{r}$ and classical computations shows that for every $\lambda\geq 0$ and $t,s\geq 0$:

### v2

@@ 4.1 Wasserstein gradient flows in the case $\lambda>0$
Let the assumptions of Lemma 3.1 hold and let $d\geq 3$. Consider some initialization $\mu_{0}\in\mathcal{P}(\Omega)$ and, for some $N\geq 0$, denote by $\hat{\mu}_{0}\coloneqq N^{-1}\sum_{i=1}^{N}\delta_{\omega_{i}}$ the empirical measure where $\{\omega_{i}\}_{1\leq i\leq N}$ are i.i.d. samples of $\mu_{0}$. For $\lambda>0$, let $(\mu^{\lambda}_{t})_{t\geq 0}$ and $(\hat{\mu}_{t}^{\lambda})_{t\geq 0}$ be the gradient flow of $\mathcal{L}^{\lambda}_{f}$ starting from $\mu_{0}$ and $\hat{\mu}_{0}$ respectively. Then there exists a constant $A=A(d,\Omega)$ s.t. for every $t\geq 0$ and every $\varepsilon>0$:
$\mathbb{P}\left(\mathcal{W}_{1}(\hat{\mu}^{\lambda}_{t},\mu^{\lambda}_{t})\geq\varepsilon\right)\leq\frac{A}{\varepsilon}N^{-1/d}e^{Ct}\,,$
where $C=C(f^{*},\lambda)$ is the constant in Theorem 1.
@@
for some constant $A=A(d,\Omega)>0$ depending on the dimension and on the domain $\Omega$. Then using the contraction rate in Theorem 1 we have a constant $C=C(f^{*},\lambda)>0$ such that for every $t\geq 0$:
@@
The result then follows by applying Markov’s inequality. ∎
@@ 4.2 Wasserstein gradient flows in the case $\lambda=0$ and ultra-fast diffusions
$\partial_{t}\mu_{t}=-\|\bar{\nu}\|^{r}_{\mathrm{TV}}\mathrm{div}\left(\mu_{t}\nabla\left(\frac{\bar{\mu}}{\mu_{t}}\right)^{r}\right).$ (45)
@@
$\partial_{t}\mu_{t}=-r\|\bar{\nu}\|^{r}_{\mathrm{TV}}\mathrm{div}\left(\bar{\mu}\nabla\left(\frac{\bar{\mu}}{\mu_{t}}\right)^{r-1}\right)\,.$
@@
Definition 2 (Gradient flow of $\mathcal{L}^{0}_{r}$ (Def. 1.1 in [iacobelli2019weighted])).
Let $\mu_{0}\in\mathcal{P}(\Omega)$ admit a density $\mu_{0}\in L^{r+2}(\Omega)$. We say $(\mu_{t})_{t\geq 0}$ is a weak solution of Eq. 45 or a gradient flow for $\mathcal{L}^{0}_{r}$ starting from $\mu_{0}$ if it is a narrowly continuous curve in $\mathcal{P}(\Omega)$ with $\lim_{t\to 0^{+}}\mu_{t}=\mu_{0}$, s.t.
@@
$\left(\frac{\mu_{t}}{\bar{\mu}}\right)^{r-1}\in L^{2}_{\mathrm{loc}}((0,\infty),H^{1}(\Omega)),\quad\frac{\bar{\mu}}{\mu_{t}}\in L^{2}_{\mathrm{loc}}((0,\infty),H^{1}(\Omega))\,.$
@@
In [iacobelli2019weighted], the authors establish the existence and uniqueness of gradient flows for the functional $\mathcal{L}^{0}_{r}$. More precisely, they show that, under appropriate assumptions on the initialization $\mu_{0}$ and on the target $\bar{\mu}$, the iterates of the proximal scheme in Eq. 38 converge towards a curve $(\mu_{t})_{t\geq 0}$ that is a gradient flow of the functional $\mathcal{L}^{0}_{r}$ in the sense of Definition 2.
Theorem 2 ([iacobelli2019weighted, Thm. 1.2]).
Assume $\mu_{0}$ and $\bar{\mu}$ are absolutely continuous and have bounded log-densities. Then there exists a unique weak solution of Eq. 45 starting from $\mu_{0}$ in the sense of Definition 2.
@@
Theorem 3 ([iacobelli2019weighted, Thm. 1.4]).
Assume $\mu_{0}$ and $\bar{\mu}$ are absolutely continuous and have bounded log-densities. For $\mu_{0}\in\mathcal{P}(\Omega)$, let $(\mu_{t})_{t\geq 0}$ be a weak solution of Eq. 45 starting from $\mu_{0}$ in the sense of Definition 2. Then in fact the log-density of $\mu_{t}$ is bounded uniformly over $t\geq 0$ and there exists a constant $C=C(\Omega,\bar{\mu},\mu_{0})>0$ s.t. for every $t\geq 0$:
@@
For completeness, we give here some of the key arguments of the proof of the above Theorem 3 in the case where $r=2$. In this case, we have for every $\mu\in\mathcal{P}(\Omega)$:
@@
where $\chi^{2}$ is the chi-square divergence. The following Lemma 4.3 establishes the desired linear convergence rate for the proximal scheme defined in Eq. 38 with the loss $\mathcal{L}^{0}_{2}$. The result in continuous time then follows from the lower semicontinuity of the $\chi^{2}$-divergence as the curve $(\mu_{t})_{t\geq 0}$ is obtained by taking the limit of the discrete process $(\mu^{\tau}_{k})_{k\geq 0}$ when the discretization time $\tau$ tends to zero.
From a technical perspective, the proof of Lemma 4.3 relies on a Poincaré inequality satisfied by $\mu_{t}$. It is indeed well-known that such inequality controls the convergence rate of Fokker-Planck equations towards their stationary distribution in $\chi^{2}$-distance [pavliotis2014stochastic, Thm. 4.4]. This can for example be used to prove the convergence of sampling algorithms such as Langevin Monte Carlo [chewi2024analysis, chewi2020svgd]. In our case, the ultra-fast diffusion Eq. 45 is to be interpreted as a Wasserstein gradient flow for $\mathcal{L}^{0}_{2}$, which, by the above Eq. 47, is the reverse $\chi^{2}$-divergence between $\mu_{t}$ and $\bar{\mu}$ and the convergence rate is controlled by the Poincaré constant of $\mu_{t}$. This rate may thus a priori evolve and vanish during training but, crucially, [iacobelli2019weighted, Lem. 2.4] shows that it is here a property of solutions to the ultra-fast diffusion equation that the log-density ratio $\|\log\left(\frac{\bar{\mu}}{\mu_{t}}\right)\|_{\infty}$ decreases with time. As a consequence, it is sufficient to assume that the log-density is bounded at initialization to obtain a control over the Poincaré constant of $\mu_{t}$, for $t\geq 0$, by a classical perturbation argument [ane2000inegalites, Thm. 3.4.1].
@@
Remark 4.2 (Dependence of the convergence rate w.r.t. the dimension).
It follows from the proof that the convergence rate $C$ in Lemma 4.3 scales linearly with $C_{P}(\Omega)^{-1}$ where $C_{P}(\Omega)$ is the Poincaré constant of the domain $\Omega$. For bounded, Lipschitz and convex domains of $\mathbb{R}^{n}$ or for the flat torus $\mathbb{T}^{n}$, this constant is in particular independent of the dimension $n$ [payne1960optimal]. Therefore, the correspondence between the training of neural networks in the two-timescale regime and solutions to ultra-fast diffusions points towards the fact that gradient methods, with suitable hyperparameter scaling, are amenable to efficient feature learning in the training of neural networks, without suffering from the curse of dimensionality [donoho2000high]. Note however that the convergence rate $C$ in Lemma 4.3 is exponentially bad in the log-density ratio $\|\log(\bar{\mu}/\mu_{0})\|_{\infty}$. In particular the convergence rate does not hold in case the teacher feature distribution is supported on a finite number of atoms.
@@ 5 Convergence of gradient flow
The main purpose of this work is to study in what extent the gradient flow dynamics defined Definitions 1 and 2 allow recovering the teacher feature distribution $\bar{\mu}$ associated to the observed signal $Y$ in Assumption 1. Whereas Theorem 3 shows convergence of the gradient flow of $\mathcal{L}^{0}_{f}$, that is solutions to the ultra-fast diffusion Eq. 45, such dynamics are hardly numerically tractable in practice due to the absence of the regularization parameter $\lambda$. For this reason we are interested here in the asymptotic behavior of the gradient flow of $\mathcal{L}^{\lambda}_{f}$ in the case where $\lambda>0$. A difficulty is that, in the case $\lambda=0$, the proof of Theorem 3 relies on the implicit behavior of Eq. 45 which preserves the density of solutions. Such a behavior is a priori not expected to hold when $\lambda>0$. As a consequence, the results in this section hold under supplementary regularity assumptions on the solutions to Eq. 41.
@@ 5.1 Algebraic convergence rate
At fixed $\lambda>0$, we are able to obtain convergence towards the minimizer $\bar{\mu}^{\lambda}_{f}$ of $\mathcal{L}^{\lambda}_{f}$ under mild regularity assumptions on solutions to the gradient flow Eq. 41. Specifically, for a probability measure $\mu\in\mathcal{P}(\Omega)$ and a function $h\in\mathcal{C}^{1}$ we define the weighted Sobolev seminorm of $h$ as:
@@
Then for a measure $\nu\in\mathcal{M}(\Omega)$ s.t. $\int_{\Omega}\mathrm{d}\nu=0$ the negative weighted Sobolev seminorm $\|\nu\|_{\dot{H}^{-1}(\mu)}$ is defined by duality with $\dot{H}^{1}(\mu)$:
@@
The following Theorem 4 states convergence of $\mathcal{L}^{\lambda}_{f}$ towards 00 with an algebraic convergence rate provided $\|\mu_{t}-\frac{\bar{\nu}}{\bar{m}}\|_{\dot{H}^{-1}(\mu_{t})}$ stays bounded along the gradient flow. As discussed below, since the domain $\Omega$ is compact, this assumption is satisfied for example when both distribution have bounded log-densities. The arguments are similar to the one presented in [glaser2021kale], where the authors consider an infimal convolution between a kernel discrepancy and the $\mathrm{KL}$-divergence. Importantly, the obtained convergence rate depends on the bound on $\|\mu_{t}-\frac{\bar{\nu}}{\bar{m}}\|_{\dot{H}^{-1}(\mu_{t})}$ but is independent of $\lambda>0$.
Theorem 4.
@@
Note that it follows from the assumptions on $f$ and $\bar{\nu}$ that $\inf\mathcal{L}^{\lambda}_{f}=0$ and that this infimum is attained only for $\mu=\bar{\nu}/\bar{m}$. Thus, the last statement on the convergence of $\mu_{t}$ follows from the convergence of $\mathcal{L}^{\lambda}_{f}(\mu_{t})$ to 00 and from the lower semicontinuity of $\mathcal{L}^{\lambda}_{f}$ (see Section 3).
To obtain the convergence rate, consider $\mu\in\mathcal{P}(\Omega)$ and note that by Eq. 27 we have:
$\mathcal{L}^{\lambda}_{f}(\mu)=\max_{\alpha}\mathcal{K}(\alpha,\mu)\,,$
where for every $\alpha\in L^{2}(\rho)$ we defined:
@@
with $f^{*}$ the Legendre transform of $f$. Let us denote by $\alpha^{\lambda}=\alpha^{\lambda}_{f}[\mu]$ the maximizer of $\mathcal{K}(\alpha,\mu)$. Then, using the convexity of $f^{*}$, we have for every $\omega\in\Omega$:
@@
Also by assumption $\partial f(\bar{m})=0$ and hence by properties of the Legendre transform $\partial f^{*}(0)=\bar{m}$. Also $f^{*}(0)=-f(\bar{m})=0$ and after integrating w.r.t. $\bar{\nu}$ :
@@
Let us comment on the assumptions of Theorem 4. The second assumption specifically, is automatically satisfied in case $\bar{\nu}$ has bounded density and $\mu_{t}$ has bounded log-density, uniformly over $t\geq 0$. Indeed, for $\mu\in\mathcal{P}(\Omega)$ having a lower-bounded log-density, we have that the weighted Sobolev seminorm .$\|.\|_{\dot{H}^{1}(\mu)}$ is lower-bounded by the classical Sobolev seminorm .$\|.\|_{\dot{H}^{1}(\pi)}$, where we recall that $\pi$ is the (normalized) Lebesgue measure over $\Omega$. Precisely, if $\pi\ll\mu$ and $\mathrm{d}\pi/\mathrm{d}\mu\leq C_{1}$ then for every $f\in\mathcal{C}^{1}$:
@@
In this case, the weighted negative Sobolev seminorm .$\|.\|_{\dot{H}^{-1}(\mu)}$ is upper-bounded by the seminorm .$\|.\|_{\dot{H}^{-1}(\pi)}$ and for every $\nu\in\mathcal{M}(\Omega)$ with $\int_{\Omega}\mathrm{d}\nu=0$ we have:
@@
Moreover, this last quantity can be estimated by the Wasserstein distance. Indeed, for probability measures having bounded log-densities, the Wasserstein distance $\mathcal{W}_{2}$ is equivalent to the negative Sobolev seminorm .$\|.\|_{\dot{H}^{-1}(\pi)}$. If $\mu,\nu\in\mathcal{P}(\Omega)$ are such that $\frac{\mathrm{d}\mu}{\mathrm{d}\pi},\frac{\mathrm{d}\nu}{\mathrm{d}\pi}\leq C_{2}$ for some constant $C_{2}>0$ we have [santambrogio2015optimal, Lem. 5.33 and Thm. 5.34]:
@@
Finally, the Wasserstein distance $\mathcal{W}_{2}(\mu,\nu)$ is always bounded by $\mathrm{diam}(\Omega)$ which is finite, hence ensuring the second assumption of Theorem 4 is satisfied.
@@
The algebraic convergence rate stated in the above Theorem 4 in the case $\lambda>0$ stands in contrast with the faster linear convergence stated in Theorem 3 in the case $\lambda=0$. For this reason, we are interested in comparing the gradient flow dynamics with and without regularization.
Below we assume $f(t)=|t|^{r}/(r-1)$ for some $r>1$ and Theorem 5 shows local uniform in time convergence of gradient flows of $\mathcal{L}^{\lambda}_{r}$ to gradient flows of $\mathcal{L}^{0}_{r}$, i.e. solutions to the ultra-fast diffusion equation Eq. 44, when the regularization strength $\lambda$ vanishes. To obtain such a result we assume regularity on the density ratio $\frac{\mathrm{d}\bar{\nu}}{\mathrm{d}\mu^{\lambda}_{t}}$. Namely, we assume that the Legendre-conjugate $\partial f(\frac{\mathrm{d}\bar{\nu}}{\mathrm{d}\mu^{\lambda}_{t}})$ stays bounded in the RKHS $\mathcal{H}$, defined as the image of the convolution operator $\Phi^{\top}:L^{2}(\rho)\to\mathcal{C}^{0}(\Omega)$ (Eq. 61). Using classical results from the theory of inverse problems, such a source condition ensures the dual variable $\alpha\in L^{2}(\rho)$ stays uniformly bounded for $\lambda>0$ (Lemma 5.1). Therefore, provided $\mathcal{H}$ is sufficiently regular, such a regularity assumption ensures compactness of the Wasserstein gradient $\nabla\mathcal{L}^{\lambda}_{r}[\mu_{t}]=\nabla f^{*}(\Phi^{\top}\alpha^{\lambda}_{t})$ in $\mathcal{C}^{1}$ and allows passing to the limit in Eq. 42 to obtain Eq. 46.
Theorem 5.
Let Assumption 1 hold with $\bar{\nu}$ a positive measure with bounded log-density, consider the regularization function $f(t)=|t|^{r}/(r-1)$ for some $r>1$ and let the assumptions of Theorem 1 and Theorem 2 be satisfied. Consider some initialization $\mu_{0}\in\mathcal{P}(\Omega)$ s.t. $\mu_{0}$ has bounded log-density. For $\lambda\geq 0$, let $(\mu^{\lambda}_{t})_{t\geq 0}$ be the gradient flow of $\mathcal{L}^{\lambda}_{r}$ starting from $\mu_{0}$ in the sense of Definition 1 (when $\lambda>0$) and Definition 2 (when $\lambda=0$). Moreover, for $\mathcal{H}$ defined by Eq. 61, assume $\mathcal{H}$ is compactly embedded in $\mathcal{C}^{1}(\Omega)$ and $\partial f(\frac{\mathrm{d}\bar{\nu}}{\mathrm{d}\mu^{\lambda}_{t}})$ is bounded in $\mathcal{H}$, locally uniformly over $t\geq 0$ and uniformly over $\lambda>0$. Then for any $T\geq 0$:
@@
For $\lambda>0$, the curves $(\mu_{t})^{\lambda}$ are gradient flows for the functionals $\mathcal{L}^{\lambda}_{r}$ and classical computations show that for every $t,s\geq 0$: